Question 4

Company XYZ had the following information during the year:

1. Sales revenue of $500,000.

2. Cost of goods sold of $38,000.

3. Operating expenses of $85,200.

4. Unrealized holding gain from available-for-sale securities of $12,500.

5. Foreign currency translation losses of $2,210.

6. Cash holdings of $101,150.

7. Beginning retained earnings of $21,000.

8. Beginning accumulated other comprehensive income of $32,000.

Compute (1) net income, (2) other comprehensive income, (3) comprehensive income, and (4) accumulated other comprehensive income.
Answer #1

1) Net income

Sales revenue 500000
Cost of goods sold -38000
Operating expense -85200
Net income 376800

2) Other comprehensive income

Unrealized holding gain from available-for-sale securities 12500
Foreign currency translation losses -2210
Other comprehensive income 10290

3) Comprehensive income = Net income+Other comprehensive income = 376800+10290 = 387090

4) Accumulated other comprehensive income = 32000+10290 = 42290